How much of an 80% solution has to be added to a 20% solution to make 960 mL of a 50% solution?
Answer by jorel1380(3719)

.80p+.20(960-p)=960(.50)
.8p+192-.2p=480
.6p=288
p=288/.6=480
960-p=480
You need 480 ml of 80% solution to mix with 480 ml of 20% solution to make your mixture..